# Temporal Evolution of the Third Interstellar Comet 3I/ATLAS: Spin, Color, Spectra and Dust Activity††thanks: based on observations made with the Southern African Large Telescope (SALT)

### v1

@@ A.1.1 Faulkes Telescopes North and South
We used the 2.0-m Faulkes Telescope North (FTN; MPC code F65), located at Haleakalā Observatory in Maui, Hawai‘i (USA), and the 2.0-m Faulkes Telescope South (FTS; MPC code E10), located at Siding Spring Observatory in New South Wales (Australia), to obtain photometry in the SDSS , , , and bands using the MuSCAT3 and MuSCAT4 instruments, respectively. MuSCAT3 is equipped with four 1k 1k CCDs (Hamamatsu Photonics), each with a pixel scale of 0.27″ and a field of view of 7.4 7.4 arcmin2, while MuSCAT4 features four 2k 2k CCDs, with a pixel scale of 0.38″ and a field of view of 13 13 arcmin2 (Narita et al. 2020; Fukui et al. 2022). Both instruments enable simultaneous multi-band imaging via dichroic beam splitters. The detectors exhibit low readout noise (3–5 e-), and the photometric data were processed using the MuSCAT pipeline, which includes standard calibrations (bias, dark, and flat-field corrections) and aperture photometry optimized for each band.
@@ A.1.4 Lesedi
Multi-filter photometric observations were collected with the f/8, 1-meter Lesedi telescope (Worters et al. 2016), located at the South African Astronomical Observatory’s (SAAO) observing site in the Northern Cape, South Africa (GPS coordinates: 32°22′47.2′′ S, 20°48′38.4′′ E; MPC code: M28). The instrument used was Mookodi (Erasmus et al. 2024), which in imaging mode provides a field of view of 10 arcmin 10 arcmin and is equipped with a full SDSS filter set (). The telescope operates in a fully robotic mode each night as part of the SAAO’s “Intelligent Observatory” (IO) initiative (Potter et al. 2024; Erasmus et al. 2024, 2025), and data were collected under the rapid-response near-Earth asteroid follow-up program currently running at the SAAO (PIs: T.S Ngwane and N. Erasmus).
@@ A.1.5 Skalnaté Pleso Observatory
Photometric material at the Skalnaté Pleso Observatory (MPC code: 056) for the interstellar comet 3I/ATLAS was acquired using a 0.61-m f/4.3 Newtonian telescope equipped with an SBIG ST-10XME CCD camera and Johnson-Cousins and photometric filters. Observations were conducted using binning, yielding a resolution of 1.07 arcsec per pixel. Standard reduction procedures were applied, including bias subtraction, dark frame subtraction, and flat-field correction. Tycho Tracker software (Parrott 2020) was used as the main tool for deriving magnitudes and quantities. Stellar magnitudes for standard stars were taken from the ATLAS catalog, an all-sky reference catalog containing approximately one billion stars down to a limiting magnitude of 19. The catalog integrates data from Pan-STARRS DR1, ATLAS Pathfinder, ATLAS re-flattened APASS, SkyMapper DR1, APASS DR9, Tycho-2, and the Yale Bright Star Catalog. Developed by Tonry et al. (2018), it provides a robust foundation for photometric calibration (Kostov & Bonev 2018). Interstellar comet 3I/ATLAS was observed on five nights between July 3 and July 23.
@@ A.2.2 Nordic Optical Telescope
Appendix B Photometry and dust parameters derivation
@@
The comet magnitudes in various filters on 02, 07, and 15 July 2025, measured within an aperture ranging from 10,300 to 11,600 km in radius centered on the optocentre, are listed in Table 1. Standard star magnitudes were taken from the APASS catalog (Henden et al. 2016), with an average photometric uncertainty of 0.02 mag.
@@
Appendix C Determination of the spin state and lightcurve amplitude
@@
The photometric measurements were reduced by applying light-time correction to the observation epochs, as well as distance corrections for the varying heliocentric and geocentric distances. A phase correction was applied using the H–G photometric system (Bowell et al. 1989), assuming a standard slope parameter of . To determine the best-fitting rotation period, we performed a Fourier analysis using a second-order harmonic function (see Fig 6).

### v2

@@ A.1.1 Faulkes Telescopes North and South
We used the 2.0-m Faulkes Telescope North (FTN; MPC code F65), located at Haleakalā Observatory in Maui, Hawai‘i (USA), and the 2.0-m Faulkes Telescope South (FTS; MPC code E10), located at Siding Spring Observatory in New South Wales (Australia), to obtain photometry in the SDSS , , , and bands using the MuSCAT3 and MuSCAT4 instruments, respectively. MuSCAT3 is equipped with four 1k 1k CCDs (Hamamatsu Photonics), each with a pixel scale of 0.27″ and a field of view of 7.4 7.4 arcmin2, while MuSCAT4 features four 2k 2k CCDs, with a pixel scale of 0.38″ and a field of view of 13 13 arcmin2 (Narita et al., 2020; Fukui et al., 2022). Both instruments enable simultaneous multi-band imaging via dichroic beam splitters. The detectors exhibit low readout noise (3–5 e-), and the photometric data were processed using the MuSCAT pipeline, which includes standard calibrations (bias, dark, and flat-field corrections) and aperture photometry optimized for each band.
@@ A.1.4 Lesedi
Multi-filter photometric observations were collected with the f/8, 1-meter Lesedi telescope (Worters et al., 2016), located at the South African Astronomical Observatory’s (SAAO) observing site in the Northern Cape, South Africa (GPS coordinates: 32°22′47.2′′ S, 20°48′38.4′′ E; MPC code: M28). The instrument used was Mookodi (Erasmus et al., 2024), which in imaging mode provides a field of view of 10 arcmin 10 arcmin and is equipped with a full SDSS filter set (). The telescope operates in a fully robotic mode each night as part of the SAAO’s “Intelligent Observatory” (IO) initiative (Potter et al., 2024; Erasmus et al., 2024, 2025), and data were collected under the rapid-response near-Earth asteroid follow-up program currently running at the SAAO (PIs: T.S Ngwane and N. Erasmus).
@@ A.1.5 Skalnaté Pleso Observatory
Photometric material at the Skalnaté Pleso Observatory (MPC code: 056) for the interstellar comet 3I/ATLAS was acquired using a 0.61-m f/4.3 Newtonian telescope equipped with an SBIG ST-10XME CCD camera and Johnson-Cousins and photometric filters. Observations were conducted using binning, yielding a resolution of 1.07 arcsec per pixel. Standard reduction procedures were applied, including bias subtraction, dark frame subtraction, and flat-field correction. Tycho Tracker software (Parrott, 2020) was used as the main tool for deriving magnitudes and quantities. Stellar magnitudes for standard stars were taken from the ATLAS catalog, an all-sky reference catalog containing approximately one billion stars down to a limiting magnitude of 19. The catalog integrates data from Pan-STARRS DR1, ATLAS Pathfinder, ATLAS re-flattened APASS, SkyMapper DR1, APASS DR9, Tycho-2, and the Yale Bright Star Catalog. Developed by Tonry et al. (2018), it provides a robust foundation for photometric calibration (Kostov & Bonev, 2018). Interstellar comet 3I/ATLAS was observed on five nights between July 3 and July 23.
@@ A.2.2 Nordic Optical Telescope
Appendix B Morphology of the coma
The absence of distinct morphological structures in comet 3I is consistent with its currently low level of activity. Long-period and dynamically new comets can display enhanced activity (Meech & Svoren, 2004), with noticeable coma structures even at large heliocentric distances (Jewitt & Luu, 2019; Ivanova et al., 2019, 2021, 2023). However, such behavior strongly depends on the intrinsic properties of the nucleus, including its volatile inventory and the spatial distribution of active areas. For comparison, comet 2I/Borisov exhibited much stronger activity, characterized by an elongated coma, jet-like outflows, and a well-developed tail (Manzini et al., 2020). The limited morphology observed in comet 3I therefore reflects both its early activation stage and its particular physical characteristics.
Appendix C Photometry and dust parameters derivation
@@
The comet magnitudes in various filters on 02, 07, and 15 July 2025, measured within an aperture ranging from 10,300 to 11,600 km in radius centered on the optocentre, are listed in Table 1. Standard star magnitudes were taken from the APASS catalog (Henden et al., 2016), with an average photometric uncertainty of 0.02 mag.
@@
Appendix D Determination of the spin state and lightcurve amplitude
@@
The photometric measurements were reduced by applying light-time correction to the observation epochs, as well as distance corrections for the varying heliocentric and geocentric distances. A phase correction was applied using the H–G photometric system (Bowell et al., 1989), assuming a standard slope parameter of . To determine the best-fitting rotation period, we performed a Fourier analysis using a second-order harmonic function (see Fig 5).
In some lightcurve segments—particularly from F65 and Z23—the rotational modulation was not clearly detected, showing a relatively flat curve clustering near the average magnitude. This likely results from observing conditions (e.g., seeing, telescope resolution) and intrinsic comet activity, which enhances the coma and dilutes nucleus variability. The extensive Z23 dataset was probably acquired under such conditions, partially or entirely suppressing the rotational signature. Inclusion of these flat segments reduced the rotational contrast and introduced ambiguity in the Fourier analysis, likely contributing to alias solutions.